Question 1

Question 1 Not yet answered Marked out of 4.00 Flag question If the number density of electrons in a wire is n 1.00 x 1028 m-3, at what drift velocity must they travel through a wire of diameter d 0.560 mm to deliver a current of 7.20 A? Give your answer in mm s1 to 3 significant figures Take T 3.14, and the magnitude of the charge on an electron e 1.60 x 10-19 C. (4 marks) Drift velocity mm s1

Question 2:

Question 2 Not yet answered Marked out of4.00 P Flag question A wire is placed at right angles (θ = 90.0°) to the lines of magnetic flux between the poles of a horseshoe magnet. The length of the wire in the field is 57.0 mm and the field strength is 0.710 T. If the current in the wire is 3.40 A (a) Calculate the force on the wire, giving your answer in newtons (N) to 3 significant figures. (2 mark) Force when 8-90.0° (b) If the wire is rotated so that it is 30.0° to the magnetic field, calculate the new force, giving your answer in newtons (N) to 3 significant figures. (2 mark) Force when 0 30.0

Question 3:

Question 3 Not yet answered Marked out of4.00 P Flag question Calculate the total power delivered by a three-phase electricity supply, where the r.m.s. input voltage Vin-350 V. Power is delivered along a wire which has resistance R1-0.170 Ω Give your answer in megawatts (Mw), to 3 significant figures. (4 mark) Power delivered MW

Question 4:

Question 4 Not yet answered Marked out of 8.00 Flag question A d.c. motor runs on a voltage of 2.70 V. The motor speed is 3750 rpm, the current is 0.480 A, and the torque is 1.90 x 103 N m. Taking T 3.14, and giving your answers in watts (W) to 3 significant figures, calculate: (a) the input power to the motor. (2 marks) nput power (b) the output power of the motor. (4 marks) Output power (c) Giving your answer as a percentage to 3 significant figures, calculate the efficiency of the motor. (2 Efficiency

Question 5:

Question 5 Not yet answered Matkad out o 10.00 Fng questin (a) Calculate the tums ratio of a step-down trangformer that transforms 8.80 KV (em.s) delivered by a power line to a 240V(r.m.s.) mains supply. Give your answwer to 3 significant Sigures.(1 mark) Turns ratio: (b) If an applance with a 26.0 hesting element is plugged into the 240V (m.s)mains supply what wil be the power supplied to the appliance? Give your answer in kilowatts (kw) to 3 significant figures. (4 marks) Power supplied kw to provide 12.0V(rm& to a 62.0 W light bulb. A mains supply of 120 V (m.s) is used across the primary coil What turns ratio is required? Give your anawer as an exact number. (1 mark) Turng ratio What cufment will be taken from the supply (primary cuiment.assuming no transformer losses? Give your answer in amps (A) to 3 significant figures. (4 marks)

Question 6:

Question 6 Not yet answered Marked out of 800 Flag question (a) Ignoring atmospheric pressure, calculate the pressure at a point in the Atlantic Ocean, 8314 m below the surface.Assume a constant density of seawater of 1025 kg m-3 Take g- 9.81 m s2 Give your answer in megapascals (MPa) to 3 significant figures. (4 marks) Pressure: MPa (b) Calculate the force exerted by the water on the circular viewing window of a submersible operating a depth of 4707 m below the surface, if the radius of the window is 0.400 m. Ignore the effect of atmospheric pressure and the air pressure inside the submersible and assume a constant density of seawater of 1025 kg m3 Take T 3.14 and g-9.81 ms2 Give your answer in meganewtons (MN) to 3 significant figures (4 marks) Force

Question 7:

Question 7 Not yet answered Marked out of 800 Flag question An incompressible fluid flows in a horizontal circular pipe of non-constant diameter. You may assume the pressure of the fluid to be constant. The fluid has dynamic viscosity 1.002 x 10- Pas, and density 998 kg m3 At point 1 the pipe diameter is 87.0 mm and the fluid velocity is 1.60 ms1 At point 2 the pipe diameter is 45.0 mm (a) Calculate the fluid velocity at point 2. Take π-3.14, and give your answer in m s-1 to 3 significant figures. (4 marks) Velocity at point 2: m s-1 (b) Calculate the Reynolds number at point 2. Give your answer to 3 significant figures. (4 marks) Hint: take care with units Reynolds number

Question 8:

Question 8 Not yet answered Marked out of 10.00Flag question Water flows through an open channel of hydraulic diameter 1.50 m, whose wetted perimeter is 6.00 m. The slope is 0.0110, and the Chézy coefficient is 5.00 m5 s1 (a) Calculate the hydraulic radius. Give your answer in m, to 3 significant figures. (2 mark) Hydraulic radius: (b) Calculate the velocity of the water flow. Give your answer in ms1, to 3 significant figures. (4 marks) Flow velocity m s1 (c) Calculate the rate of discharge of the flowing water. Give your answer in cubic metres per second (m3 s1), to 3 significant figures. (4 marks) Rate of discharge m3 s-1

Please number each question. Thank you
Answer #1

Question No 01:-

Drift velocity is given by

Vd = I/(neA) = 7.20A/(1*10^28m^-3*1.6*10^-19C*π(0.560*10^-3/2)^2)

= 4.567*10^-3 m/s

Question no :02

Magnetic force is given by F = Bil*sin©

when © = 90°

F = 0.710*3.40*57*10^-3*sin90° = 0.1376 N

when sin© = 30

F = 0.710*3.40*57*10^-3sin30° = 0.0688 N

Question 03:-

Power delivered = V^2/R = 350^2/0.170 = 0.1721 MW